What is the size of the force on an electron when the electron is in a uniform electric field that has an intensity of 8.7 x 10 7 N/C
1
Expert's answer
2020-02-25T10:44:43-0500

"F=E \cdot e=8.7 \cdot 10^7 \cdot 1.6 \cdot 10^{-19}=1.392 \cdot10^{-11}N"